Circular No. 7415

Central Bureau for Astronomical Telegrams INTERNATIONAL ASTRONOMICAL UNION

Mailstop 18, Smithsonian Astrophysical Observatory, Cambridge, MA 02138, U.S.A.

IAUSUBS@CFA.HARVARD.EDU or FAX 617-495-7231 (subscriptions) BMARSDEN@CFA.HARVARD.EDU or DGREEN@CFA.HARVARD.EDU (science)

        URL http://cfa-www.harvard.edu/iau/cbat.html ISSN 0081-0304

Phone 617-495-7244/7440/7444 (for emergency use only)

VARIABLE STAR IN FIELD OF NGC 3432 M. Papenkova and W. D. Li, University of California at Berkeley, on

behalf of the Lick Observatory Supernova Search (cf. IAUC 6627, 7126), report the discovery with the 0.8-m Katzman Automatic Imaging Telescope (KAIT) of a variable star (thought not to be a supernova) in the field of NGC 3432. The new object was discovered and confirmed on unfiltered images taken on May 3.2 (mag ∼ 17.4) and 4.2 UT (mag ∼ 18.2), respectively. The new object is located at α = 10^{h}52^{m}41^{s}.40, δ = +36^{o}40′08′′.5 (equinox 2000.0), which is ∼ 123′′ east and 180′′ north of the diffuse nucleus of NGC 3432 and superimposed on an H II region (or spiral arm) of the galaxy. An unfiltered image of the field on Apr. 29.2 (limiting mag ∼ 19.0) showed nothing at the position of the new object.

SUPERNOVA 2000cb IN IC 1158 J. Y. Wei, Y. L. Qiu, J. S. Deng, and J. Y. Hu, Beijing Astronomical

Observatory (BAO), on behalf of the BAO supernova survey, report their detection with the 0.6-m reflector of SN 2000cb (cf. IAUC 7410) on unfiltered images taken on Apr. 29.8, 30.8, and May 3.7 UT, all of which showed the supernova at mag ∼ 16.3. Nothing was visible at the position of SN 2000cb on a similar image obtained on Apr. 7.8.

COMET C/1999 J2 (SKIFF) T. Fukushima, T. Nakajima, and J. Watanabe, National Astronomical

Observatory of Japan (NAOJ), report that I-band CCD images of comet C/1999 J2 taken with the NAOJ 0.50-m telescope show a dust antitail at the large heliocentric distance of 7 AU (the comet having passed perihelion on Apr. 5): Apr. 26.687 UT, 2′.60 in p.a. 12^{o}; May 2.576, 3′.31 in p.a. 18^{o}. Coma diameters were ∼ 0′.4 on both dates, and m_{1} = 15.6 in V and 14.9 in I on Apr. 29. The earth will cross the comet’s orbital plane on May 10.4.

V463 SCUTI = NOVA SCUTI 2000 Visual magnitude estimates of this nova (cf. IAUC 7382) by A. Pearce,

Nedlands, W. Australia: Mar. 17.883 UT, 11.5; 23.849, 13.0; 25.810, 12.9; 26.858, 13.3; 27.846, 13.6; 28.837, 13.9; 29.835, 13.9; 30.783, 14.0; Apr. 1.851, 14.2; 2.849, 14.2; 5.840, 14.2; 7.840, 14.3; 8.788, 14.5; 9.843, 14.5.

2000 May 5  \begin{matrix}c \\ ◯ \end{matrix} Copyright 2000 CBAT  Daniel W. E. Green